### v1

@@ 3.2 Description of the simulation database
[…]
@@ 3.3 Model constants for the hexagonal sphere pack
[…]
@@
Figure 4: Forchheimer equation (15) with coefficients (3.3) and direct numerical simulation data for stationary flow through a hexagonal sphere pack.
@@ 3.4 Implementation of the models
The correctness of the scheme (3.4.1) and its convergence with the time step size is assessed using the method of manufactured solutions. As a test case, we consider the velocity $\left\langle u\right\rangle_{\mathrm{s}}=j_{0}\,\frac{t^{2}}{2}\,\theta(t)$ with the Heaviside function $\theta(t)$. The forcing that is necessary to find this velocity as a solution to (13) can be obtained as
@@
We simulated (3.4.1) with the forcing defined above for a hexagonal sphere pack. Figure 4(a) shows the relative error of $\left\langle u\right\rangle_{\mathrm{s}}$ with respect to the analytic solution at $\frac{\nu T}{d^{2}}=5$. We observe first order convergence which means that the singularity in the integral kernel is treated with sufficient accuracy.
@@ 4.1 Comparison of model errors
The error of the model predictions with respect to the direct numerical simulation results is quantified using the cycle-averaged $L^{2}$ error (37); we further decompose this error into an amplitude (41) and a phase contribution (42) using a Fourier series. Table 3 shows the amplitude and phase contribution to the $L^{2}$ error of the different model predictions with respect to the direct numerical simulation dataset. Note that the linear models were not applied to the nonlinear cases since they cannot account for the nonlinear drag and thus would produce large errors at high Reynolds numbers. In almost all cases, the main source of error is the amplitude error. In contrast, the phase error is significant only in four cases in which it reaches about half the size of the amplitude error.
@@
[…]
@@ 4.2 Analysis of prediction errors
[…]

### v2

@@ 3.2 Description of the simulation database
and therefore depends on two dimensionless parameters: The Hagen number $\mbox{{Hg}}=\left|g_{x}\right|d^{3}/(\rho\nu^{2})$ determines the amplitude and the Womersley number $\mbox{{Wo}}=\sqrt{\Omega\,d^{2}/\nu}$ determines the frequency of the macroscopic pressure gradient. The Reynolds number was defined as
$\mbox{{Re}}=\limsup\limits_{t\to\infty}\frac{\left|\left\langle\boldsymbol{u}\right\rangle_{\mathrm{s}}\right|d}{\nu}$ (31)
and was obtained as a result of the simulations. The simulations were performed at the three Womersley numbers $\mbox{{Wo}}=10$, $31.62$, and $100$ that correspond to the low, medium and high frequency regime in linear flow, respectively.222 The medium frequency regime in linear flow through a hexagonal sphere pack occurs around a Womersley number $\mbox{{Wo}}_{0}=\sqrt{\epsilon\,d^{2}/(\alpha_{\infty}\,K)}=30.5$, where $\Omega_{0}=\epsilon\,\nu/(\alpha_{\infty}\,K)$ is the transition frequency (Pride et al, 1993). The geometric constants are given in table 2. The Hagen numbers were set such that the simulations cover linear flow, laminar nonlinear flow, and transitional and turbulence-like flow at each Womersley number. The flow was started from rest and simulated until a recurrent behaviour could be observed in the superficial velocity. The difference in the superficial velocity in the last simulated cycle and the preceding cycle was less than approximately $0.1\%$ of the maximum superficial velocity in the last simulated cycle for the laminar cases and about $2\%$ for the transitional and turbulence-like cases. The simulation parameters can be found in table 1.
@@ 3.3 Model constants for the hexagonal sphere pack
The coefficients of the Forchheimer equation (15) are determined based on the simulation results in (Sakai and Manhart, 2020; Unglehrt and Manhart, 2023a). The linear Forchheimer coefficient $a$ is chosen such that the Forchheimer equation approaches Darcy’s law for $\mbox{{Re}}\to 0$. Then, the nonlinear Forchheimer coefficient $b$ is determined from a least squares fit to the ratio of the macroscopic pressure gradient and the superficial velocity. We obtained the following coefficients:
@@
Figure 4: Forchheimer equation (15) with coefficients (32) and direct numerical simulation data for stationary flow through a hexagonal sphere pack.
@@ 3.4 Implementation of the models
The correctness of the scheme (33) and its convergence with the time step size is assessed using the method of manufactured solutions. As a test case, we consider the velocity $\left\langle u\right\rangle_{\mathrm{s}}=j_{0}\,\frac{t^{2}}{2}\,\theta(t)$ with the Heaviside function $\theta(t)$. The forcing that is necessary to find this velocity as a solution to (13) can be obtained as
@@
We simulated (33) with the forcing defined above for a hexagonal sphere pack. Figure 5(a) shows the relative error of $\left\langle u\right\rangle_{\mathrm{s}}$ with respect to the analytic solution at $\frac{\nu T}{d^{2}}=5$. We observe first order convergence which means that the singularity in the integral kernel is treated with sufficient accuracy.
@@ 4.1 Comparison of model errors
The error of the model predictions with respect to the direct numerical simulation results is quantified using the cycle-averaged $L^{2}$ error (37); we further decompose this error into an amplitude (41) and a phase contribution (42) using a Fourier series. Table 3 shows the amplitude and phase contribution to the $L^{2}$ error of the different model predictions with respect to the direct numerical simulation dataset. Note that the linear models were not applied to the nonlinear cases since they cannot account for the nonlinear drag and thus would produce large errors at high Reynolds numbers. Both the amplitude error and the phase error make significant contributions to the overall error of the models. The two components are weakly correlated, with the phase error being on average about $57\%$ of the amplitude error, but the importance of the two components varies considerably between models and simulation cases.
@@
We find that the extended dynamic permeability model based on the model of Pride et al (1993) has very small errors in linear flow and for flow at low frequencies. The reason for this is that the model equation (23) reverts to the model of Pride et al (1993) as $\mbox{{Re}}\to 0$ and to the steady Forchheimer equation as $\Omega/\Omega_{0}\to 0$. The prediction errors are largest for nonlinear flow at high frequencies. Interestingly, the prediction error for the case HF7 ($\mbox{{Re}}=3580$) is smaller than for the case HF6 ($\mbox{{Re}}=1080$). The model of Turo and Umnova (2013) performs similar to the extended (Pride et al, 1993) model at high frequencies; however, the prediction errors in the medium and low frequency regime are very large. The reason for this is the excessive damping in the linear regime that was discussed in section 2.3. The unsteady Forchheimer equation with the acceleration coefficient based on the static viscous tortuosity $\alpha_{0}$ has relatively small errors at low frequencies, moderate errors at medium frequencies and very large errors at high frequencies. The largest errors can be observed in nonlinear flow at high frequencies. In contrast, the unsteady Forchheimer equation with the acceleration coefficient based on the high-frequency limit of the dynamic tortuosity $\alpha_{\infty}$ has comparably small errors at low frequencies, very large errors in linear flow at medium and high frequencies and small errors for flow at high Reynolds numbers. This somewhat surprising behaviour of the unsteady Forchheimer equation with $\alpha_{\infty}$ will be investigated in the following section.
@@ 4.2 Analysis of prediction errors
In this section, we aim to explain the different error behaviours described above. We first investigate the discrepancies in the linear regime, where the model of Turo and Umnova (2013) (equation 18) and the unsteady Forchheimer equation (16) have large errors at low and high frequencies, respectively. We then proceed to the nonlinear high frequency regime where large discrepancies between the models can be observed and the flow state is characterised by the interaction between strong accelerations and turbulence. Therefore, we take a closer look at the predictions of the models for the linear case HF1 and for the turbulent case HF6 at $\Omega/\Omega_{0}=10.7$ ($\mbox{{Wo}}=100$). Figure 7 shows the predictions of the different models in comparison to the superficial velocity from the simulations.